# Exciton fractional Chern insulators in moiré heterostructures

### v1

@@ III Abelian states at half filling
We now consider the realistic long-range exciton–exciton interactions introduced in Eq. (3) and take into account the impact of the small but finite kinetic energy. The results of exact diagonalization and the subsequent analysis are remarkably similar to those obtained for contact interactions, cf. Fig. 2(e)-(h). The fact that the energy gap between ground and excited many-body states ($\sim 10$ meV) persists for the considered systems ranging from $N_{\text{s}}=10$ up to $N_{\text{s}}=20$ sites clearly reflects that such phases are robust (see Appendix C.1). Thus, our numerical calculations demonstrate that exciton FCIs are stable in the considered moiré heterostructure.
@@
Interestingly, the calculated gap of $\sim 10$ meV is two times larger than in the case of hole FCIs in the similar system of twisted bilayer MoTe2 [55]. In principle, this suggests a higher stability of exciton FCIs compared to their electron/hole counterparts, whose gap has been experimentally estimated to be 20 K [56]. The potentially larger gap of exciton FCIs might be a result of the stronger repulsive interaction dominated by $V^{\text{e-e}}+V^{\text{h-h}}$ in the short range ($V^{\text{e-h}}$ is weaker since the charges are located in separate layers) instead of just $V^{\text{e-e}}$ or $V^{\text{h-h}}$. A larger gap for bosons can also be expected from a pseudopotential perspective, where the interaction for bosons and fermions is dominated by the zeroth, $v_{0}$ ($\delta(\mathbf{r)}$ component of the interaction) and first $v_{1}$ ($-a^{2}_{\text{m}}\nabla^{2}\delta(\mathbf{r)}$ component) pseudopotentials, respectively, with the bosonic interaction being stronger than the fermionic one (generally $v_{n}>v_{n+1}$ in lowest Landau level-like bands) [49, 8]. We note, though, that a reliable quantitative estimation of the gap is difficult due to limitations such as the finite system size, the use of a pure Coulomb potential instead of the thin-film Rytova-Keldysh potential [21], and uncertainties in parameters such as the dielectric constant.
IV Non-Abelian states at filling one
@@
After confirming the presence of Abelian exciton FCIs, we now seek exotic phases whose elementary excitations obey non-Abelian anyon statistics. In this context, the most straightforward and promising candidate is the Moore–Read state [57], which emerges as the exact zero-energy ground state at filling $\nu=1$ for bosons in the lowest Landau level with artificial three-body contact interactions [58] and which was predicted to appear in rotating Bose-Einstein condensates [59, 60]. In the following, we show that this phase is stable in the realistic conditions considered here, i.e. in the moiré band and assuming long-range interactions.
@@
Figure 3: Moore–Read states at filling one. Many-body spectrum (10 lowest energies) and PES at $\nu=1$ with long-range interactions for (a)-(b) 13 and (c)-(d) 14 particles. The single and threefold quasi-degenerate ground states (red dots in panels (a) and (c)) for an odd and even number of sites are located at the momenta expected for Moore–Read states. The red line in panels (b) and (d) indicates the entanglement gap respecting the quasi-hole counting rules of the Moore–Read state (416 and 518 quasihole excitations for 13 and 14 particles with $N_{A}=3$). PES for larger $N_{A}$ are shown in Appendix C.2.
@@
First, we note that the most prominent feature of the Moore–Read phase is the dependence of the ground state degeneracy on the parity of the particle number. Here, the thin-torus exclusion principle [61] states that, at most, two particles can occupy two consecutive orbitals [62, 54], resulting in three possible Fock-space configurations, $111111\cdots$, $202020\cdots$, and $020202\cdots$. As a result of periodic boundary conditions, only the first configuration is allowed for an odd number of particles—resulting in a single ground state—, while for an even number of particles all three configurations are allowed and the ground state is threefold degenerate.
@@
In Fig. 3(a),(c), we show the calculated many-body spectra for systems with 13 and 14 excitons. The single and three lowest states in the case of odd and even number of excitons, respectively, are located at the momenta expected from the thin-torus exclusion rule for Moore–Read states. We note that the spread of the ground state energies and the small gap with respect to excited states is a common feature of finite-size systems. We expect the states to become exactly degenerate in the thermodynamic limit and note that the ground and excited states remain well separated upon threading magnetic flux (see Appendix C.2). Moreover, the calculation of many-body Chern number yields $\mathcal{C}_{\text{avg}}=1$ for each state and reveals a uniform Berry curvature distribution, indicating that the Moore–Read phase is robust.
@@
In order to unambiguously determine the nature of the phase beyond the parity dependence of the ground-state degeneracy, we calculate the PES. For both $N_{\text{x}}=13$ and $N_{\text{x}}=14$ systems, the PES is characterized by a large gap, below which the number of states exactly matches the counting of quasi-hole excitations allowed by the Moore–Read exclusion rule. Furthermore, the PES gap persists across various system sizes and different $N_{A}$ (see Appendix C.2). Taken together, the ground-state degeneracy and momenta, the many-body Chern number, and the PES for systems with different particle-number parity constitute convincing evidence for the presence of a robust non-Abelian Moore–Read phase with exciton constituents at filling $\nu=1$ in the considered moiré heterostructure.
V Conclusion and outlook
@@
We have explored the emergence of strongly-correlated topological phases arising from a many-body bosonic system of long-lived moiré excitons in a flat Chern band. In particular, by combining exact diagonalization with many-body diagnosis tools, we have demonstrated the existence of robust Abelian and non-Abelian exciton fractional Chern insulators in a realistic model of a moiré heterostructure at filling $\nu=\frac{1}{2}$ and $\nu=1$, respectively. Our work introduces a versatile and accessible platform for investigating exciton FCIs and opens a new avenue towards the realization of non-Abelian anyons.
Importantly, our findings provide a specific guideline for the experimental realization of exciton FCIs in twisted bilayer WSe2 stacked on monolayer MoSe2. We note that, despite the exciton’s neutral net charge, there are methods which enable exciton transport that could be utilized to detect these phases. The most promising route is likely to be counterflow transport measurements [63, 64]. Other methods for studying transport of interlayer excitons involve creating spatial gradients in the out-of-plane electric field, dielectric environment, or strain profiles [65, 66]. Future studies should also attempt to identify the optical fingerprints of these phases.
@@
Apart from prompting experimental efforts in a specific system, our work motivates the search for exciton FCIs in other moiré structures. Furthermore, while we have only focused on two specific phases at fillings $\nu=\frac{1}{2}$ and $\nu=1$, other highly nontrivial phases remain unexplored. Besides other bosonic Laughlin and hierarchy states, the pursuit of additional non-Abelian phases such as the Moore–Read state at $\nu=\frac{1}{3}$ and the Read–Rezayi state [67] at $\nu=\frac{3}{2}$ is particularly interesting—especially so for the latter as it hosts Fibonacci anyons which hold large promise for topological quantum computing.
@@ Acknowledgements.
Appendix A Moiré exciton model
A.1 Exciton basis
Electrons and holes in each layer can be described in an effective mass approach at the K valley with the Hamiltonian $H_{\text{e-h}}=H_{\text{e},0}+H_{\text{h},0}+H_{\text{e-h, int}}$, where $H_{\text{e},0}=\sum_{\mathbf{k}}E^{\text{e}}_{\mathbf{k}}e^{\dagger}_{\mathbf{k}}e^{\phantom{\dagger}}_{\mathbf{k}}$, $H_{\text{h},0}=\sum_{l\mathbf{k}}E^{\text{h}}_{\mathbf{k}}h^{\dagger}_{l\mathbf{k}}h^{\phantom{\dagger}}_{l\mathbf{k}}$, and $H_{\text{e-h, int}}=-\sum_{l\mathbf{kk^{\prime}q}}V^{\text{e-h}}_{l\mathbf{q}}e^{\dagger}_{\mathbf{k+q}}h^{\dagger}_{l\mathbf{k^{\prime}-q}}h^{\phantom{\dagger}}_{l\mathbf{k}^{\prime}}e^{\phantom{\dagger}}_{\mathbf{k}}$. Here, $E^{\text{e}}_{\mathbf{k}}=\hbar^{2}\mathbf{k}^{2}/2m_{\text{e}}$, $E^{\text{h}}_{\mathbf{k}}=E^{\text{gap}}_{l}+\hbar^{2}(\mathbf{k}-\mathbf{K}_{l})^{2}/2m_{\text{h}}$, $\mathbf{K}_{l}$ is the K-point for holes at the layer $l=1,2$ and $V^{\text{e-h}}_{l\mathbf{q}}=\frac{e_{0}^{2}}{2A\epsilon\epsilon_{0}|\mathbf{q}|}\mathrm{e}^{-|\mathbf{q}|dl}$ is the electron-hole interaction potential with the interlayer distance $d$, the dielectric constant $\epsilon$, and the system area $A$. Considering that all electrons and holes are paired into tightly bound excitons, the system can be described in terms of the 1s exciton state, |$\ket{X_{l\mathbf{Q}}}=\sum_{\mathbf{k}}\phi_{l\mathbf{k}}e^{\dagger}_{\mathbf{k}_{\text{e}}}h^{\dagger}_{\mathbf{k}_{\text{h}}}\ket{0}$, where |$\ket{0}$ is the semiconductor ground state, $\phi_{l\mathbf{k}}$ is the exciton wave function with the relative momentum $\mathbf{k}=\alpha_{\text{h}}\mathbf{k}_{\text{e}}-\alpha_{\text{e}}(\mathbf{k}_{\text{h}}-\mathbf{K}_{l})$, $\alpha_{\text{e/h}}=m_{\text{e/h}}/(m_{\text{h}}+m_{\text{e}})$, and $\mathbf{Q}=\mathbf{k}_{\text{h}}-\mathbf{K}_{l}+\mathbf{k}_{\text{e}}$ is the center-of-mass momentum. The exciton dispersion $E^{\text{x}}_{l\mathbf{Q}}=E^{\text{gap}}_{l}+E^{\text{x}}_{l}+\frac{\hbar^{2}\mathbf{Q}^{2}}{2M_{\text{x}}}$ is determined by the mass $M_{\text{x}}=m_{\text{h}}+m_{\text{e}}$ and the offset energy $E^{\text{x}}_{l}$ that fulfills the eigenvalue problem $H_{\text{e-h}}\ket{X_{l\mathbf{Q}}}=E^{\text{x}}_{l\mathbf{Q}}\ket{X_{l\mathbf{Q}}}$, which corresponds to the Wannier equation
$\frac{\hbar^{2}\mathbf{k}^{2}}{2\mu_{\text{x}}}\phi_{l\mathbf{k}}-\sum_{\mathbf{q}}V^{\text{e-h}}_{l\mathbf{q}}\phi_{l,\mathbf{k+q}}=E^{\text{x}}_{l}\phi_{l\mathbf{k}},$ (4)
where $\mu_{\text{x}}=m_{\text{e}}m_{\text{h}}/M_{\text{x}}$ is the reduced mass. The exciton wave function is well approximated by a hydrogen ansatz, which in real space reads $\phi_{l}(\mathbf{r})\propto\mathrm{exp}(-|\mathbf{r}|/a_{l})$. Taking $m_{\text{h}}=0.43m_{0},m_{\text{e}}=0.8m_{0}$, $d=0.67$ nm and $\epsilon=3.8$[35] we obtain the exciton binding energies $E^{\text{b}}_{1}=230\ \text{meV},E^{\text{b}}_{2}=140\ \text{meV}$ and Bohr radii $a_{1}=1.3\ \text{nm},a_{2}=1.8\ \text{nm}$ via variational minimization, in close agreement with the exact numerical solution of Eq. (4).
A.2 Exciton–exciton interaction

### v2

@@ Abstract
We now consider realistic long-range exciton–exciton interactions (see End Matter) and take into account the impact of the small but finite kinetic energy. The long-range interaction potential contains electron–electron, hole–hole, and electron–hole interaction mechanisms and is governed by the direct component, which at long distances behaves as the interaction between two dipoles with lengths $dl$ and $dl^{\prime}$, i.e. $V_{ll^{\prime}}(\mathbf{r})\sim(dl)(dl^{\prime})/|\mathbf{r}|^{3}$. Formally, the interaction potential also contains exchange terms, but typically these play a minor role in the context of interlayer excitons [59, 60, 61, 62]. Considering these realistic exciton–exciton interactions, the results of exact diagonalization and the subsequent analysis are remarkably similar to those obtained for contact interactions, cf. Fig. 2(e)-(h). The fact that the energy gap between ground and excited many-body states ($\sim 10$ meV) persists for the considered systems ranging from $N_{\text{s}}=10$ up to $N_{\text{s}}=20$ sites clearly reflects that such phases are robust (see SM). Thus, our numerical calculations demonstrate that exciton FCIs are stable in the tWSe2–MoSe2 moiré structure.
Furthermore, we show that the many-body energy gap remains large ($\gtrsim 10$ meV) for a broad range of twist angles $\theta$ and energy offsets $\Delta$, cf. Fig.3(a). In particular, the Laughlin-like FCI phase seems to be stable as long as the single-particle exciton band is topological (see SM), and a sizable gap persists even at large twist angles ($\theta\sim 3^{\circ}$) where the band is not flat anymore, cf. Fig. 3(b). Our calculations indicate that the $\nu=\frac{1}{2}$ exciton FCI phase is most robust at twist angles $\theta\in[2^{\circ},2.5^{\circ}]$ and offsets $\Delta\in[-5\ \text{meV},0\ \text{meV}]$, where the many-body and PES gaps are largest and the energy spread between the two ground states is smallest (see SM).
@@
Interestingly, the calculated gap of $\sim 10$ meV is two times larger than in the case of hole FCIs in the similar system of twisted bilayer MoTe2 [63]. In principle, this suggests a higher stability of exciton FCIs compared to their electron/hole counterparts, whose gap has been experimentally estimated to be 20 K [64]. The potentially larger gap of exciton FCIs might be a result of the stronger repulsive interaction dominated by $V^{\text{e-e}}+V^{\text{h-h}}$ in the short range ($V^{\text{e-h}}$ is weaker since the charges are located in separate layers) instead of just $V^{\text{e-e}}$ or $V^{\text{h-h}}$. A larger gap for bosons can also be expected from a pseudopotential perspective, where the interaction for bosons and fermions is dominated by the zeroth $v_{0}$ and first $v_{1}$ pseudopotentials, respectively, where generally $v_{n}>v_{n+1}$ in lowest Landau level-like bands [53, 8]. We note, though, that a reliable quantitative estimation of the gap is difficult due to limitations such as the finite system size, the use of a pure Coulomb potential [21], and uncertainties in parameters such as the dielectric constant.
Figure 3: Stability of Laughlin states in the $(\theta,\Delta)$ parameter space, where $\theta$ is the twist angle and $\Delta$ is the energy offset between the two interlayer exciton resonances. (a) Energy gap in the many-body spectrum between the Laughlin ground states and excited states at filling $\nu=\frac{1}{2}$ in a system of $N_{\text{s}}=12$ sites with long-range interaction. (b) Bandwidth of the lowest single-particle exciton band.
@@
Non-Abelian states at filling one—After confirming the presence of Abelian exciton FCIs, we now seek exotic phases whose elementary excitations obey non-Abelian anyon statistics. In this context, the most straightforward and promising candidate is the MR state [65]. This state can be theoretically understood as the exact zero-energy ground state at filling $\nu=1$ for bosons in the lowest Landau level with artificial three-body contact interactions [66] and it has been predicted to appear in rotating Bose-Einstein condensates [67, 68]. In principle, MR states can be stabilized by long-range two-body interactions [69], which are at least more favourable than (two-body) short-range pseudopotentials for achieving the required pairing. In addition, the bosonic quantum statistics might further facilitate the pairing and, thus, the stabilization of the typically elusive MR states. In the following, we show numerical evidence indicating that this phase is stable in the realistic conditions considered here, i.e. in the moiré exciton band and assuming long-range interactions.
@@
Figure 4: Moore–Read states at filling one. Many-body spectrum (10 lowest energies) and PES at $\nu=1$ with long-range interactions for (a)-(b) 13 and (c)-(d) 14 particles. The single and threefold quasi-degenerate ground states (red dots in panels (a) and (c)) for an odd and even particle number of sites are located at the momenta expected for MR states. The red line in panels (b) and (d) indicates the entanglement gap respecting the quasi-hole counting rules of the MR state (416 and 518 quasihole excitations for 13 and 14 particles with $N_{A}=3$).
@@
First, we note that the most prominent feature of the MR phase is the dependence of the ground state degeneracy on the parity of the particle number. Here, the thin-torus exclusion principle [70] states that, at most, two particles can occupy two consecutive orbitals [71, 58], resulting in three possible Fock-space configurations, $111111\cdots$, $202020\cdots$, and $020202\cdots$. As a result of periodic boundary conditions, only the first configuration is allowed for an odd number of particles—resulting in a single ground state—, while for an even number of particles all three configurations are allowed and the ground state is threefold degenerate.
@@
In Fig. 4(a),(c), we show the calculated many-body spectra for systems with 13 and 14 excitons. The single and three lowest states in the case of odd and even number of excitons, respectively, are located at the expected momenta for MR states. We note that the spread of the ground state energies and the small gap with respect to excited states is a common feature of finite-size systems [13, 14, 15, 16, 19]. We expect the states to become exactly degenerate in the thermodynamic limit, and we note that the gap persists upon threading magnetic flux and for different system sizes (see SM). Moreover, the calculation of many-body Chern number yields $\mathcal{C}_{\text{avg}}=1$ for each state and reveals a uniform Berry curvature distribution, indicating that the MR phase is robust.
@@
In order to unambiguously determine the nature of the phase beyond the parity dependence of the ground-state degeneracy, we calculate the PES. For both $N_{\text{x}}=13$ and $N_{\text{x}}=14$ systems, the PES is characterized by a large gap, below which the number of states exactly matches the counting of quasi-hole excitations allowed by the MR exclusion rule. Furthermore, the PES gap persists across various system sizes and different $N_{A}$ (see SM). All together, these results constitute convincing evidence for the presence of a robust non-Abelian MR phase with exciton constituents at filling $\nu=1$ in the tWSe2–MoSe2 moiré structure.
@@
Conclusion and outlook—We have explored the emergence of strongly-correlated topological phases arising from a bosonic system of long-lived moiré excitons in a flat Chern band. In particular, by combining exact diagonalization with many-body diagnosis tools, we have demonstrated the existence of robust Abelian and non-Abelian exciton fractional Chern insulators in a realistic model of the tWSe2–MoSe2 moiré heterostructure at filling $\nu=\frac{1}{2}$ and $\nu=1$, respectively. Our work establishes a versatile and accessible platform for investigating exciton FCIs and opens a promising avenue towards the realization of non-Abelian anyons.
Importantly, our findings provide a specific guideline for the experimental realization of exciton FCIs in tWSe2–MoSe2. Concretely, we predict that the exciton FCI remains stable across a wide range of experimental parameters and is particularly robust at $\theta\in[2^{\circ},2.5^{\circ}]$ and $\Delta\in[-5\ \text{meV},0\ \text{meV}]$. The experimental know-how for fabricating moiré heterostructures and generating long-lived interlayer excitons is very mature [27, 24, 29, 33]. The experimental challenges will likely lie on the detection of the topological order and perhaps on the control of the exciton density. We note that, despite the exciton’s neutral net charge, there are methods which enable exciton transport that could be utilized to detect these phases—the most promising are counterflow transport measurements [72, 73], while other methods might involve spatial gradients in the out-of-plane electric field, dielectric environment, or strain profiles [74, 75]. Future studies could also attempt to identify the optical fingerprints of these phases. The experimental control of the density of interlayer excitons with microsecond lifetimes is steadily improving [76]. We also note that introducing hBN spacers between electron and hole layers further enhances the lifetime and also allows for the direct injection of interlayer excitons via electrical gates [43]. Importantly, even if the exciton density is not accurately fixed to the relevant filling factor, the particles can, in principle, re-arrange themselves into FCI puddles (pinned by disorder) with the appropriate density, similar to the presumable origin of Hall resistance plateaus as a function of gate voltage in fermionic moiré FCIs.
@@
Apart from prompting experimental efforts in tWSe2–MoSe2, our work motivates the search for exciton FCIs in other moiré structures. Furthermore, while we have only focused on two specific phases at $\nu=\frac{1}{2}$ and $\nu=1$, other phases remain unexplored. Besides other bosonic Laughlin and hierarchy states, the pursuit of additional non-Abelian phases such as the Moore–Read state at $\nu=\frac{1}{3}$ and the Read–Rezayi state [77] at $\nu=\frac{3}{2}$ is particularly interesting—especially so for the latter as it hosts Fibonacci anyons which hold large promise for topological quantum computing. Our work serves as a catalyst for the exploration of exotic physics including unprecedented phenomena unique to many-boson systems.
@@ Acknowledgements.
Appendix A Exciton–exciton interaction